# Unveiling the Potential of Multimodal Retrieval Augmented Generation with Planning

### v1

@@ Abstract.
Multimodal Retrieval Augmented Generation (MRAG) systems, while promising for enhancing Multimodal Large Language Models (MLLMs), often rely on rigid, single-step retrieval methods. This limitation hinders their ability to effectively address real-world scenarios that demand adaptive information acquisition and query refinement. To overcome this, we introduce the novel task of Multimodal Retrieval Augmented Generation Planning (MRAG Planning), focusing on optimizing MLLM performance while minimizing computational overhead. We present CogPlanner, a versatile framework inspired by human cognitive processes. CogPlanner iteratively refines queries and selects retrieval strategies, enabling both parallel and sequential modeling approaches. To rigorously evaluate MRAG Planning, we introduce CogBench, a new benchmark specifically designed for this task. CogBench facilitates the integration of lightweight CogPlanner with resource-efficient MLLMs. Our experimental findings demonstrate that CogPlanner surpasses existing MRAG baselines, achieving significant improvements in both accuracy and efficiency with minimal computational overhead.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Retrieval Augmented Generation (RAG) has proven to be a significant advancement in enhancing the generation quality of large language models (LLMs) (Izacard and Grave, 2020; Lewis et al., 2020). While RAG frameworks have primarily focused on textual information, the burgeoning need for more sophisticated systems demands the incorporation of multimodal data, including images and videos (Riedler and Langer, 2024). This evolution has given rise to Multimodal Retrieval Augmented Generation (MRAG), which aims to enable multimodal large language models (MLLMs) to access and leverage multimodal external knowledge sources, thereby mitigating hallucinations and enhancing the system reliability (Zhao et al., 2023a). However, challenges lie in the profound intricacies of real-world human interactions where the users tend to provide low-resolution screenshots (Menges et al., 2018) and succinct queries with heavy reliance on anaphoric references (Spink et al., 2001). This behavior necessitates accurate information-seeking and integration mechanisms to construct comprehensive responses. Existing MRAG frameworks generally adhere to a rigid pipeline characterized by a predetermined retrieval action, either exclusively textual or exclusively visual (Chen et al., 2022; Zhu et al., 2024), which manifests several critical limitations:
@@
  • Blind Information Acquisition: As demonstrated in recent studies (Shi et al., 2023; Asai et al., 2023), the compulsive retrieval mechanism without proper consideration of necessity or relevance can introduce irrelevant contextual information that undermines the MLLM’s capability for accurate responses. Moreover, it neglects the inherent capabilities of MLLMs to reason and process multimodal data, rendering the retrieval step redundant or counterproductive.
@@
  • Inadequate Query Formation: The visual incompleteness, textual ambiguity, and conciseness create a fundamental impediment that fails to retrieve pertinent information. Furthermore, the current single-step MRAG methodologies prove inadequate for addressing multi-hop queries that require multi-step reasoning process (Kil et al., 2024).
@@
In response to these real-world limitations, we introduce a novel task, termed Multimodal Retrieval Augmented Generation Planning (MRAG Planning), which determines an optimal procedure for multimodal query processing, enabling judicious information seeking and integration. The objective is to maximize the capacity of the underlying MLLMs to generate a comprehensive and accurate response while minimizing the computational cost overhead to the MRAG system. To this end, the MRAG Planning task is formally decomposed into two interrelated sub-tasks: (1) Information acquisition which discerns the truly necessary information required by the MLLMs and devises the appropriate retrieval strategies accordingly. (2) Query reformulation which involves decomposing complex, multi-hop queries into manageable atomic sub-queries, and refining them for clarity and informativeness. Figure 1 illustrates a multi-hop reasoning example where the first step is to identify the game name AstroBot through an image search. The initial query is then decomposed into two sub-queries, each focusing on the sales data for the two respective games. Subsequent web searches are conducted to retrieve sales data and generate the final response. This example underscores the crucial need for dynamic planning procedures in MRAG systems where both order and selection of retrieval methods, and query reformulation are determined tailored to the specific characteristics of the multimodal query.
@@
To address the challenges of MRAG Planning, we propose a novel framework, CogPlanner  [^1] We name our framework as CogPlanner because of the inherent cognitive process of humans., inspired by human cognitive processes. Just as humans synthesize and gather multimodal information to address complex queries, adapting to their cognitive abilities and knowledge base, CogPlanner emulates this behavior by dynamically determining a planning procedure in coordination with downstream MLLMs. CogPlanner operates through two core actions: query reformulation and retrieval action selection. Query reformulation involves breaking down a complex query into related sub-queries or refining the queries. The retrieval strategy, in turn, encompasses image search, text search, or none. When sufficient information is gathered, the framework refrains from further retrieval, culminating in the generation of a final response. This iterative procedure reflects an adaptive chain of actions tailored to the specific needs of each multimodal query. For instance, the optimal planning strategy in Figure 1 follows (Query Refinement, Image Search) $\rightarrow$ (Query Decomposition, Text Search). CogPlanner supports two distinct modeling approaches: parallel and sequential modeling, each differs in the order of the query reformulation and retrieval action selection actions.
@@
In conjunction with the introduction of the CogPlanner framework, we present CogBench, a comprehensive dataset tailored to the decision-making processes central to the MRAG Planning task. CogBench consists of over 7,000 data samples, with a high-quality test set of 401 samples. The creation of CogBench involves two steps: (1) Query Collection, wherein we enrich with collected screenshots and multi-hop reasoning queries demanding long-form answers, and (2) MRAG Planning and Generation, wherein we apply CogPlanner with GPT-4o (Achiam et al., 2023) and human annotations. In addition to its essential role in evaluation, the development of CogBench facilitates the design of specialized fine-tuning strategies aimed at bolstering the decision-making capabilities of smaller, resource-efficient MLLMs. By utilizing the CogBench training set, we achieve lightweight integration of the Qwen2-VL-7B-Instruct (Wang et al., 2024) model as the planning expert in CogPlanner. This integration, referred to as Qwen2-7b-VL-Cog, maintains its resource-efficient characteristics while enabling effective performance within the MRAG Planning context. In summary, the contributions of this work are as follows:
@@
  • We develop CogBench benchmark, tailored to the MRAG Planning task. It supports performance evaluation and facilitates finetuning strategies to enable the lightweight integration of resource-efficient MLLMs with CogPlanner. Experimental results demonstrate that CogPlanner achieves more than 15% improvements over various MRAG approaches while incurring less than 10% additional costs with Qwen-7b-VL.
@@ 2.2. Multimodal Retrieval Augmented Generation
RAG frameworks have demonstrated considerable success in various real-world applications (Li et al., 2022). However, their reliance on textual information presents a significant limitation, as it precludes the incorporation of crucial knowledge embedded within other modalities, such as images and videos. Multimodal Retrieval Augmented Generation (MRAG) seeks to address this limitation by equipping MLLMs with access to a broader spectrum of knowledge, encompassing up-to-date and domain-specific information (Zhao et al., 2023b). Empirical studies consistently demonstrate the effectiveness of MRAG systems (Zhao et al., 2023a) across various visual question answering (VQA) benchmarks (Wu et al., 2017). Recent advancements in MRAG have demonstrated notable progress. For instance, MuRAG (Chen et al., 2022) highlights how incorporating visual information retrieved from external sources significantly improves the system performance. Other prominent approaches include Plug-and-Play (Tiong et al., 2022), which transforms visual content into textual descriptions to facilitate integration with conventional text-based question-answering mechanisms. Additionally, RAMM (Yuan et al., 2023) enhances the generation process by incorporating both text-to-image retrieval and subsequent fusion of the representations for more accurate answer generation. Further innovations include Wiki-llava (Caffagni et al., 2024) and mR2AG (Zhang et al., 2024b), which enable retrieval from online knowledge bases, such as Wikipedia, via image-based queries, to provide more contextually informed responses to user queries. M2RAG (Ma et al., 2024) extends these efforts by enabling concurrent retrieval of both textual and visual elements in response to multimodal queries, allowing for more robust query understanding and generation capabilities. Additionally, benchmarks such as MRAG-Bench (Hu et al., 2024) and MMSearch (Jiang et al., 2024b) have been introduced to evaluate MRAG performance, particularly in tasks with image-to-image retrieval, addressing challenges related to incomplete or insufficient image data.
The prevailing methodologies predominantly adhere to a rigid, single-modality search paradigm. However, in authentic user scenarios, knowledge acquisition can originate from diverse sources, contingent upon the specific query and the underlying domain. To this end, we introduce the novel task of MRAG Planning. The primary goal of MRAG Planning is to systematically determine the most effective query processing strategy tailored to each multimodal query and the underlying MLLM.
@@ 3. Task Formulation
Consider a multimodal query $\mathcal{Q}_{0}=(q,v)$ where $q$ represents the textual component and $v$ represents the visual component (e.g. an image). The objective of MRAG is to retrieve pertinent information from a document collection $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{n}\}$, and generate cogent responses. Drawing parallels with human information processing, we introduce the task of MRAG Planning that interfaces intimately with the retrieval tools and downstream MLLMs in the MRAG systems, restructuring their information gathering mechanism. We formalize the MRAG system environment as a tuple $(\mathcal{G},\mathcal{I})$. Here, $\mathcal{G}$ refers to the goal conditions of assembling sufficient information to generate comprehensive and accurate responses. $\mathcal{S}$ represents the state, capturing the current set of information available, which may encompass queries and any retrieved documents. The initial state corresponds to the input multimodal query, $\mathcal{I}=\mathcal{S}_{0}=\mathcal{Q}_{0}$. The MRAG Planning task can thus be framed as a state transition function $\mathcal{F}$ that progresses from $\mathcal{I}$ toward the goal state through a chain of decisions. Formally, this transition process is defined as follows:
@@ 3.1.2. Information Aquisition
Recognizing the inherent limitations of MLLMs in terms of specific knowledge gaps, an information acquisition mechanism is imperative to supplement the MLLM’s knowledge base. We define $\mathcal{A}$ comprising three distinct retrieval operations, text search, image search, and non-search. The optimal retrieval action is determined based on the quality of the available multimodal information and the estimated utility of additional context from external knowledge sources:
@@ 4.1. Baselines
  • Fixed textual retrieval: These methods closely resemble traditional RAG frameworks. They employ textual queries to retrieve relevant documents. Subsequently, the original user query and the retrieved information are utilized as complementary inputs to generate the final answer.
  • Fixed visual retrieval: These approaches prioritize visual information, leveraging visual queries to retrieve relevant images and their associated captions. The MLLM then integrates both the original user query and the augmented retrieved images to produce the final answer.
While these fixed workflows have been effective within certain domains, they exhibit limited flexibility. The rigid, modality-specific structure hinders their ability to adapt to other contexts, or the non-search scenario where fixed search may bring extra noise. It undermines their potential for broader applicability in dynamic, real-world scenarios. In our experimental settings, we utilize these two methodologies as baseline comparisons to evaluate the efficacy of our proposed CogPlanner.
@@ 4.2. CogPlanner
We propose CogPlanner, a flexible, plug-and-play framework that mirrors the human cognitive processes when handling complex multimodal queries. Our approach is motivated that this process exhibits both adaptability and efficiency in managing multimodal information acquisition and integration. When confronted with such complex multimodal queries, humans instinctively engage in a methodical, multi-step process that involves continuous evaluation of information requirements, determination of appropriate retrieval mechanisms for the missing knowledge, and decomposition and refinement of complex queries into manageable sub-components. What distinguishes this process is its iterative nature, where each step builds upon previous understanding, ultimately converging on an information-rich and query-clear state. Crucially, each individual naturally modulates their planning based on their knowledge base and cognitive capabilities.
CogPlanner operationalizes this cognitive architecture through an iterative decision-making framework. It dynamically orchestrates a chain of actions for each query, optimizing both effectiveness and efficiency in conjunction with the downstream MLLMs. As shown in Figure 2, the multimodal query — seeking up-to-date news about Luigi’s arrest — requires three processing rounds to gather sufficient information for the current GPT knowledge base. CogPlanner centers on two critical decisions in $\mathcal{P}$: (1) query reformulation, and (2) action selection among text search, image search, and non-search, corresponding to the sub-tasks outlined in Section 3.1. We implement the core state transition function $\mathcal{F}$ through a planning expert who makes these decisions at each iteration, completing the roles of $\mathcal{F}_{IA}$ and $\mathcal{F}_{QR}$. Then, the retrieval are invoked to assess and process relevant multimodal elements.
@@ 4.2.1. Planning Expert
We employ a MLLM as the planning expert. As formalized in Equation 2, at each iterative step, the expert analyzes the current multimodal content, evaluates the information gathered thus far, and subsequently determines the most appropriate follow-up retrieval action and reformulates the query. To accomplish this, we propose two distinct modeling paradigms - parallel modeling and sequential modeling - each differs in the order in which the decision-making occurs.

### v2

# CogPlanner: Unveiling the Potential of Agentic Multimodal Retrieval Augmented Generation with Planning

@@ Abstract.
Multimodal Retrieval Augmented Generation (MRAG) systems have shown promise in enhancing the generation capabilities of multimodal large language models (MLLMs). However, existing MRAG frameworks primarily adhere to rigid, single-step retrieval strategies that fail to address real-world challenges of information acquisition and query reformulation. In this work, we introduce the task of Multimodal Retrieval Augmented Generation Planning (MRAG Planning) that aims at effective information seeking and integration while minimizing computational overhead. Specifically, we propose CogPlanner, an agentic plug-and-play framework inspired by human cognitive processes, which iteratively determines query reformulation and retrieval strategies to generate accurate and contextually relevant responses. CogPlanner supports parallel and sequential modeling paradigms. Furthermore, we introduce CogBench, a new benchmark designed to rigorously evaluate the MRAG Planning task and facilitate lightweight CogPlanner integration with resource-efficient MLLMs, such as Qwen2-VL-7B-Cog. Experimental results demonstrate that CogPlanner significantly outperforms existing MRAG baselines, offering improvements in both accuracy and efficiency with minimal additional computational costs.
@@
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3767695.3769486conference: Make sure to enter the correct conference title from your rights confirmation emai; December 7-10, 2025; Xi’an, Chinaisbn: 979-8-4007-2218-9/2025/12.ccs: Computing methodologies Natural language generationccs: Information systems Language modelsccs: Information systems Question answering11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding author.
@@ 1. Introduction
Retrieval-Augmented Generation (RAG) has been shown to significantly enhance the performance of large language models (LLMs) by grounding generation in retrieved knowledge (Izacard and Grave, 2020; Lewis et al., 2020). More recently, the emergence of agentic RAG frameworks, exemplified by web agents (Li et al., 2025; Wu et al., 2025), has highlighted the potential of autonomous reasoning and information seeking capabilities within RAG systems. Howevewr, the increasing demands of real-world applications have necessitated a natural extension of RAG beyond purely texts to encompass multimodal data (e.g., images, videos). This development has led to the advent of Multimodal Retrieval-Augmented Generation (MRAG) (Riedler and Langer, 2024), which equips multimodal large language models (MLLMs) with the ability to retrieve and exploit external multimodal knowledge sources, thereby reducing hallucinations and improving reliability (Zhao et al., 2023a). Existing MRAG frameworks generally adhere to a rigid pipeline characterized by a predetermined retrieval action, either exclusively textual or exclusively visual (Chen et al., 2022; Zhu et al., 2024), which manifests several critical limitations:
@@
  • Blind Information Acquisition: As demonstrated in recent studies (Shi et al., 2023; Asai et al., 2023), this compulsive retrieval mechanism without proper consideration of necessity or relevance can introduce irrelevant contextual information that undermines the MLLM’s capability for accurate responses. Moreover, it neglects the inherent capabilities of MLLMs to reason and process multimodal data, rendering the retrieval step redundant or counterproductive.
@@
  • Inadequate Query Formation: The visual incompleteness, textual ambiguity, and conciseness create a fundamental impediment that fails to retrieve pertinent information. Furthermore, these existing single-step MRAG methodologies prove inadequate for addressing multi-hop queries that require a multi-step reasoning process (Kil et al., 2024).
@@
In response to these real-world limitations, we propose a new task, Multimodal Retrieval-Augmented Generation Planning (MRAG Planning). The objective is to establish optimized trajectories for information seeking and integration in multimodal queries. It guies the downstream MLLMs toward more accurate and comprehensive responses, while reducing the computational overhead. This task constitutes a central component in the development of effective agentic MRAG systems. To this end, the MRAG Planning task is decomposed into two interrelated sub-tasks: (1) Information acquisition that discerns the truly necessary information required by the MLLMs and devises the appropriate retrieval strategies accordingly. (2) Query reformulation, which involves decomposing complex, multi-hop queries into manageable atomic sub-queries and refining them for clarity and informativeness. Figure 1 illustrates a multi-hop reasoning example where the first step is to identify the game name AstroBot through an image search. The initial query is then decomposed into two sub-queries, each focusing on the sales data for the two respective games. Subsequent web searches are conducted to retrieve sales data and generate the final response. This example underscores the crucial need for dynamic planning procedures in MRAG systems where both order and selection of retrieval methods, and query reformulation are determined tailored to the specific characteristics of the multimodal query.
@@
To address the challenges of MRAG Planning, we propose a novel framework, CogPlanner  [^1] We name our framework as CogPlanner because of the inherent cognitive process of humans., inspired by human cognitive processes. Just as humans synthesize and gather multimodal information to address complex queries, adapting their reasoning based on prior knowledge, CogPlanner emulates this behavior through a centralized planning expert that dynamically determines a planning procedure in coordination with downstream MLLMs. CogPlanner operates through two core operations: query reformulation and retrieval action selection. Query reformulation involves breaking down a complex query into related sub-queries or refining the queries. The retrieval strategy, in turn, encompasses image search, text search, or none. When sufficient information is gathered, the framework refrains from further retrieval, culminating in the generation of a final response. This iterative procedure reflects an adaptive chain of actions tailored to the specific needs of each multimodal query. For instance, the optimal planning strategy in Figure 1 follows (Query Refinement, Image Search) $\rightarrow$ (Query Decomposition, Text Search). CogPlanner supports two distinct modeling approaches: parallel and sequential modeling. Each differs in the order of the query reformulation and retrieval action selection.
@@
In conjunction with the introduction of the CogPlanner framework, we present CogBench, a comprehensive dataset tailored to the MRAG Planning task. CogBench consists of over 5,000 data samples, with a high-quality test set of 401 samples. In addition to its essential role in evaluation, the development of CogBench facilitates the design of specialized fine-tuning strategies aimed at bolstering the decision-making capabilities of smaller, resource-efficient MLLMs. By utilizing the CogBench training set, we achieve lightweight integration of the Qwen2-VL-7B-Instruct (Wang et al., 2024) model as the planning expert in CogPlanner. This integration, referred to as Qwen2-7B-VL-Cog, maintains its resource-efficient characteristics while enabling effective performance within the MRAG Planning context. In summary, the contributions of this work are as follows:
@@
  • We develop the CogBench benchmark, tailored to the MRAG Planning task. It supports performance evaluation and facilitates fine-tuning strategies to enable the lightweight integration of resource-efficient MLLMs with CogPlanner. Experimental results demonstrate that CogPlanner achieves more than 15% improvements over various MRAG approaches while incurring less than 10% additional costs with Qwen2-VL-7B.
@@ 2.2. Multimodal Retrieval Augmented Generation
RAG frameworks have demonstrated considerable success in various real-world applications (Li et al., 2022). However, their reliance on textual information presents a significant limitation, as it precludes the incorporation of crucial knowledge embedded within other modalities, such as images and videos. Multimodal Retrieval Augmented Generation (MRAG) seeks to address this limitation by equipping MLLMs with access to a broader spectrum of knowledge, encompassing up-to-date and domain-specific information (Zhao et al., 2023b). Empirical studies consistently demonstrate the effectiveness of MRAG systems (Zhao et al., 2023a) across various visual question answering (VQA) benchmarks (Wu et al., 2017). Recent advancements in MRAG have demonstrated notable progress. For instance, MuRAG (Chen et al., 2022) highlights how incorporating visual information retrieved from external sources significantly improves the system performance. Other prominent approaches include Plug-and-Play (Tiong et al., 2022), which transforms visual content into textual descriptions to facilitate integration with conventional text-based question-answering mechanisms. Additionally, RAMM (Yuan et al., 2023) enhances the generation process by incorporating both text-to-image retrieval and subsequent fusion of the representations for more accurate answer generation. Further innovations include Wiki-llava (Caffagni et al., 2024) and mR2AG (Zhang et al., 2024b), which enable retrieval from online knowledge bases, such as Wikipedia, via image-based queries, to provide more contextually informed responses to user queries. M2RAG (Ma et al., 2024) extends these efforts by enabling concurrent retrieval of both textual and visual elements in response to multimodal queries, allowing for more robust query understanding and generation capabilities. Additionally, benchmarks such as MRAG-Bench (Hu et al., 2024) and MMSearch (Jiang et al., 2024b) have been introduced to evaluate MRAG performance, particularly in tasks with image-to-image retrieval, addressing challenges related to incomplete or insufficient image data. The prevailing methodologies predominantly adhere to a rigid, single-modality search paradigm. However, in authentic user scenarios, knowledge acquisition can originate from diverse sources, contingent upon the specific query and the underlying domain. To this end, we introduce the novel task of MRAG Planning. The primary goal of MRAG Planning is to systematically determine the most effective query processing strategy tailored to each multimodal query and the underlying MLLMs.
@@ 3. Task Formulation
Consider a multimodal query $\mathcal{Q}_{0}=(q,v)$ where $q$ represents the textual component and $v$ represents the visual component (e.g., an image). The objective of MRAG is to retrieve pertinent information from a document collection $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{n}\}$, and generate cogent responses. Drawing parallels with human information processing, we introduce the task of MRAG Planning that interfaces intimately with the retrieval tools and downstream MLLMs in the MRAG systems, restructuring their information gathering mechanism. We formalize the MRAG system environment as a tuple $(\mathcal{G},\mathcal{I})$. Here, $\mathcal{G}$ refers to the goal conditions of assembling sufficient information to generate comprehensive and accurate responses. $\mathcal{S}$ represents the state, capturing the current set of information available, which may encompass queries and any retrieved documents. The initial state corresponds to the input multimodal query, $\mathcal{I}=\mathcal{S}_{0}=\mathcal{Q}_{0}$. The MRAG Planning task can thus be framed as a state transition function $\mathcal{F}$ that progresses from $\mathcal{I}$ toward the goal state through a chain of decisions. Formally, this transition process is defined as follows:
@@ 3.1.2. Information Aquisition
Recognizing the inherent limitations of MLLMs in terms of specific knowledge gaps, an information acquisition mechanism is imperative to supplement the MLLM’s knowledge base. We define $\mathcal{A}$ as comprising three distinct retrieval operations: text search, image search, and non-search. The optimal retrieval action is determined based on the quality of the available multimodal information and the estimated utility of additional context from external knowledge sources:
@@ 4.1. Baselines
  • Fixed textual retrieval, which closely resembles traditional RAG frameworks. They employ textual queries to retrieve relevant documents to generate the final answer.
  • Fixed visual retrieval, which prioritizes visual information, leveraging visual queries to retrieve relevant images and their associated captions. The MLLM then integrates both the original user query and the augmented retrieved images to produce the final answer.
While these fixed workflows have been effective within certain domains, they exhibit limited flexibility. The rigid, modality-specific structure hinders their ability to adapt to other contexts, or the non-search scenario, where fixed search may bring extra noise. It undermines their potential for broader applicability in dynamic, real-world scenarios.
@@ 4.2. CogPlanner
We propose CogPlanner, a flexible, plug-and-play framework that mirrors the human cognitive processes when handling complex multimodal queries. Our approach is inspired by the observation that humans demonstrate both adaptability and efficiency in acquiring and integrating multimodal information. When confronted with such queries, individuals instinctively engage in a structured, multi-step process that involves continuous assessment of information gaps, determination of appropriate retrieval strategies for the missing knowledge, and decomposition and refinement of complex queries into manageable sub-components. Crucially, this process is inherently adaptive - guided by the individual’s prior knowledge and cognitive capabilities, with the goal of converging toward a state of conceptual clarity and informational completeness.
CogPlanner operationalizes this cognitive architecture through an iterative decision-making framework. It dynamically orchestrates a chain of decisions for each query, optimizing both effectiveness and efficiency in conjunction with the downstream MLLMs. As shown in Figure 2, the multimodal query — seeking up-to-date news about Luigi’s arrest — requires three processing rounds to gather sufficient information for the current GPT knowledge base. CogPlanner centers on two critical decisions in $\mathcal{P}$: (1) query reformulation, and (2) action selection among text search, image search, and non-search, corresponding to the sub-tasks outlined in Section 3.1. We implement the core state transition function $\mathcal{F}$ through a planning expert who makes these decisions at each iteration, completing the roles of $\mathcal{F}_{IA}$ and $\mathcal{F}_{QR}$. Then, the retrieval is invoked to assess and process relevant multimodal elements.
@@ 4.2.1. Planning Expert
We employ an MLLM as the planning expert. As formalized in Equation 2, at each iterative step, the expert analyzes the current multimodal content, evaluates the information gathered thus far, and subsequently determines the most appropriate follow-up retrieval action and reformulates the query. To accomplish this, we propose two distinct modeling paradigms - parallel modeling and sequential modeling - each of which differs in the order in which the decision-making occurs.